OpenStudy (anonymous):

What is the equation of the line y= -0.25x + 1 in standard form? Use integer coefficients.

OpenStudy (anonymous):

It is standard !

OpenStudy (anonymous):

no its in slope intercept form y=mx+b its asking for standard form

OpenStudy (anonymous):

y is alone and its enough !

jigglypuff314 (jigglypuff314):

standard form of a linear equation is ax + by = c so just bring x and y together on one side

jigglypuff314 (jigglypuff314):

so for y= -0.25x + 1 add 0.25x to both sides +0.25x +0.25x 0.25x + y = 1 then multiply both sides by 4 *4 *4 *4 because it was asking for integer coefficients x + 4y = 4